Draft version 0.0

ATLAS NOTE

May 9, 2012

1 Updated search for the Standard Model Higgs boson in the decay channel 2 H → ZZ^{(∗)} → 4ℓ with 4.8 fb^{−1} of pp collisions at \sqrt{s} = 7 TeV

3 The ATLAS collaboration

4 Abstract

5 Contents

6 1 Introduction 2

7 2 Data and Simulation Samples 3

8 3 Lepton Reconstruction and Identification 5

9 4 Event Selection Optimization 6 10 4.1 Optimization for the four muons final state . . . . . . . . . . . . . . . . . . . . . . . . . 6 11 4.2 Optimization for channels with Electrons . . . . . . . . . . . . . . . . . . . . . . . . . 8

12 5 Event Selection 11

13 6 Trigger Considerations 12

14 7 Data based study of the lepton selection criteria 13 15 7.1 Efficiency for signal-like leptons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 16 7.1.1 Z → ee . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 17 7.1.2 Z → μμ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 18 7.2 Efficiency for background-like leptons . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 19 7.2.1 Z + e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 20 7.2.2 Z + μ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

21 8 Background Estimation 14 22 8.1 Data driven cross-checks of the tt̄background . . . . . . . . . . . . . . . . . . . . . . . 14 23 8.2 Estimate of the Z + μμ background . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 24 8.2.1 Efficiency estimation for muons in Jets . . . . . . . . . . . . . . . . . . . . . . 14 25 8.3 Estimate of the Z + ee background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 26 8.4 Data driven cross-checks of the QCD background . . . . . . . . . . . . . . . . . . . . . 17 27 8.5 Background shapes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

28 9 Results of Event Selection 20

29 10 Systematic Uncertainties 21

30 11 Exlusion limits and p-values 22

31 12 Conclusions 23

32 Appendices 24

33 1 Introduction 34 2 Data and Simulation Samples

35 The data are subjected to quality requirements: events recorded during periods when the relevant detector 36 components were not operating normally are rejected. The resulting integrated luminosity is 4.81 fb^{−1} , 37 4.81 fb^{−1} and 4.91 fb^{−1} for the 4μ, 2e2μ and 4e final states, respectively. 38 The H → ZZ^{(∗)} → 4ℓ signal is modelled using the powheg Monte Carlo (MC) event generator [?, ?], 39 which calculates separately the gluon-gluon and vector-boson fusion production mechanisms with matrix 40 elements up to next-to-leading order (NLO). The Higgs boson transverse momentum (p_{T}) spectrum in 41 the gluon fusion process is reweighted to match the calculation of Ref. [?], which includes quantum 42 chromodynamics (QCD) corrections up to NLO and QCD soft-gluon resummations up to next-to-next-43 to-leading logarithm (NNLL). powheg is interfaced to pythia [?] for showering and hadronization, 44 which in turn is interfaced to photos [?] for quantum electrodynamics (QED) radiative corrections in 45 the final state and to tauola [?, ?] for the simulation of τ lepton decays. pythia is used to simulate 46 the production of a Higgs boson in association with a W or a Z boson. 47 The Higgs boson production cross sections and decay branching ratios [?, ?, ?, ?], as well as their 48 uncertainties, are taken from Refs. [?, ?]. The cross sections for the gluon fusion process have been 49 calculated at next-to-leading order (NLO) in QCD [?, ?, ?], and then at next-to-next-to-leading order 50 (NNLO) [?, ?, ?]. In addition, QCD soft-gluon resummations up to NNLL are applied for the gluon fu-51 sion process [?]. The NLO electroweak (EW) corrections are applied [?, ?]. These results are compiled 52 in Refs. [?, ?, ?] assuming factorisation between QCD and EW corrections. The cross sections for the 53 vector-boson fusion process are calculated with full NLO QCD and EW corrections [?, ?, ?], and approx-54 imate NNLO QCD corrections are available [?]. The associated productions with a W or Z boson are 55 calculated at NLO [?] and at NNLO [?] in QCD, and NLO EW radiative corrections [?] are applied. The 56 uncertainty in the production cross section due to the choice of QCD scale is _{−8}^{+12} % for the gluon fusion 57 process, and ±1% for the vector-boson fusion, associated W H production, and associated ZH produc-58 tion processes [?]. The uncertainty in the production cross section due to the parton distribution function 59 (PDF) and α_{s} is ±8% for gluon-initiated process and ±4% for quark-initiated processes [?, ?, ?, ?, ?]. 60 The Higgs boson decay branching ratio to the four-lepton final state is predicted by prophecy4f [?, ?], 61 which includes the complete NLO QCD+EW corrections, interference effects between identical final-62 state fermions, and leading two-loop heavy Higgs boson corrections to the four-fermion width. Table 1 63 gives the production cross sections and branching ratios for H → ZZ^{(∗)} → 4ℓ for several Higgs boson 64 masses. 65 The cross section calculations do not take into account the width of the Higgs boson, which is im-66 plemented through a relativistic Breit-Wigner line shape applied at the event-generator level. It has been 67 suggested [?, ?, ?, ?] that effects related to off-shell Higgs boson production and interference with other 68 SM processes may become sizeable for the highest masses (m_{H} > 400 GeV) considered in this search. 69 In the absence of a full calculation, a conservative estimate of the possible size of such effects is in-70 cluded as a signal normalization systematic uncertainty following a parameterization as a function of 71 m_{H}: 150% × m_{H}^{3}[TeV], for m_{H} ≥ 300 GeV[?]. 72 The ZZ^{(∗)} continuum background is modelled using pythia. The mcfm [?, ?] prediction, including 73 both quark-antiquark annihilation and gluon fusion at QCD NLO, is used for the inclusive total cross 74 section and the shape of the invariant mass of the ZZ^{(∗)} system (m_{ZZ^{(∗)}}). The QCD scale uncertainty 75 has a ±5% effect on the expected ZZ^{(∗)} background, and the effect due to the PDF and α_{s} uncertainties 76 is ±4% (±8%) for quark-initiated (gluon-initiated) processes. An additional theoretical uncertainty of 77 ±10% on the inclusive ZZ^{(∗)} cross section is conservatively included due to the missing higher-order 78 QCD corrections for the gluon-initiated process, and a correlated uncertainty on the predicted m_{ZZ^{(∗)}} 79 spectrum is estimated by varying the gluon-initiated contribution by 100% [?]. 80 The Z +jetsproduction is modelled using alpgen [?] and is divided into two sources: Z +light jets Table 1: Higgs boson production cross sections for gluon fusion, vector-boson fusion and associated production with a W or Z boson in pp collisions at \sqrt{s} = 7TeV[?]. The quoted uncertainties correspond to the total theoretical systematic uncertainty. The production cross section for associated production with a W or Z boson is negligibly small for m_{H} > 300 GeV. The decay branching ratio for H → 4ℓ, with ℓ = e or μ, is reported in the last column [?].

m_{H} σ (gg → H) σ (qq′→ Hqq′) σ (qq̄→ W H) σ (qq̄→ ZH) BR( H → ZZ^{(∗)} → 4ℓ)
[GeV]  [pb]  [pb]  [pb]  [pb]  [10^{−3} ]
130  14.1_{−2.1}^{+2.7}  1.154_{−0.027}^{+0.032}  0.501 ± 0.020 0.278 ± 0.014  0.19
150  10.5_{−1.6}^{+2.0}  0.962_{−0.021}^{+0.028}  0.300 ± 0.012 0.171 ± 0.009  0.38
200  5.2_{−0.8}^{+0.9}  0.637_{−0.015}^{+0.022}  0.103 ± 0.005 0.061 ± 0.004  1.15
400  2.0 ± 0.3  0.162_{−0.005}^{+0.010}  −  −  1.21
600 0.33 ± 0.06  0.058_{−0.002}^{+0.005}  −  −  1.23

81 — which includes Zcc̄ in the massless c-quark approximation and Zbb̄ from parton showers — and 82 Zbb̄ using matrix element calculations that take into account the b-quark mass. The MLM [?] matching 83 scheme is used to remove any double counting of identical jets produced via the matrix element calcu-84 lation and the parton shower, but this scheme is not implemented for b-jets. Therefore, bb̄ pairs with

85 separation ΔR = \sqrt{(Δϕ)^{2}+ (Δη)^{2}} > 0.4 between the b-quarks are taken from the matrix-element calcu-86 lation, whereas for ΔR < 0.4 the parton-shower bb̄pairs are used. In this search the Z + jetsbackground 87 is normalized using control samples from data. For comparisons with simulation, the QCD NNLO 88 fewz [?, ?] and mcfm cross section calculations are used for inclusive Z boson and Zbb̄ production, 89 respectively. The tt̄background is modelled using mc@nlo [?] and is normalized to the approximate 90 NNLO cross section calculated using hathor [?]. The effect of the QCD scale uncertainty on the cross 91 section is _{−9}^{+4} %, while the effect of PDF and α_{s} uncertainties is ±7%. Both alpgen and mc@nlo are 92 interfaced to herwig [?] for parton shower hadronization and to jimmy [?] for the underlying event 93 simulation. 94 Generated events are fully simulated using the ATLAS detector simulation [?] within the GEANT4 95 framework [?]. Additional pp interactions in the same and nearby bunch crossings (pile-up) are included 96 in the simulation. The MC samples are reweighted to reproduce the observed distribution of the mean 97 number of interactions per bunch crossing in the data.

98 3 Lepton Reconstruction and Identification 99 4 Event Selection Optimization

100 Contact : F. Cerutti 101 The kinematic cuts applied at selection level have been otimized aiming at the maximal signal sen-102 sitivity. The focus has been on the low mass region, between 120 and 130 GeV where the analysis is 103 more challenging and the differences in kinematics between the signal and the reducible background are 104 larger.

105 4.1 Optimization for the four muons final state

The optimization procedure is based on a scan of the 6 main kinematic selection variables: M_{12} the opposite-charge di-muon mass coser to the Z one, M_{34} the other opposite-charge di-muons invariant mass, and the transverse momenta of the four leading muons P_{ti}with i ordered from 1 to 4 from the leading to the lowest transverse momentum muon. The position of the cut is chosen at the value that gives the maximum signal sensitivity at a given signal mass. Since the variables are correlated the otpimization procedure was performed in an iterative way looking for minima in the 6 dimentional variable space. The statistical variable used is the probability that the observed number of events, assumed to be equal to the sum of the expected background and signal events in a 6 GeV mass window around the signal mass, is as large or larger than the expected background, the expected P_{0}. This probability is computed on the basis of a 125 GeV Higgs signal and background events predicted by the MC assuming an integrated luminosity of 20 fb^{−1} at a centre-of-mass energy of 7 TeV. Two methods are used to compute the expected P_{0}: one is based on the χ^{2} asymptotic approximation with the formula described below in Equation 1. In this case the probaility is expressed as the equivalent number of standard deviations Z_{0} for a Gaussian probability density function.

Z_{0} = \sqrt{2( (s + b)ln(1 + \frac{s}{b}) − s)}  (1)

106 where s and b are the number of signal and background events, respectively. 107 The second one is based on a simple Poisson marginal probability integral with a linear extrapolation 108 between the two closest integers to the expected signal plus background number of events. Both approach 109 give consistent results in terms of the optimal cuts choice. 110 In the kinematic optimization for the four muon channel the background estimates are taken form the 111 MC (mc11c samples). This is possible since the reducible background is negligible with respect to the 112 irreducible one and therefore the optimizatin procedure is quite insensitive to its systematics. All kind 113 of backgrounds described Section 2 are taken into account in this procedure. No pileup reweighting and 114 efficiency scale factors are applied in this study. Theirs impact on the optimal cut choice is expected to 115 be negligible for the four muon channel. 116 In this optimization procedure some constraints were applied to take into account trigger require-117 ments and the fact that some of the background samples are produced with kinematic cuts at gerator 118 level so it does not make sense to explore kinematics too close to these cuts. The most sensitive variables 119 for the optimization are found to be M_{12} and P_{t4}, the lowest transverse momentum of the quadruplet. This 120 is shown in Figure 1 and Figure 2 whether the local expected Z_{0} is shown for the 4 muon channels as a 121 function of different kinematic variables. The other variables have a less strong dependency. In particular 122 P_{t1}and P_{t3}have a quite flat behavior so it has been decided to tighten the latter to 10 GeV with respect to 123 the 7 GeV of the previously pubished analysis. Also the cut on M_{34} variable has been tighten with respect 124 to the published analysis with a small increase in the expected Z_{0}. In general if a small dependency of the 125 expected Z_{0} is observed the choice has been to move the cut at the more conservetive value with respect 126 to the redicuble background contribution. The choice of the cuts based on this optimization procedure 127 are summarized in Table 2.

Figure 1: Dependency of the expected Z_{0} as described in the text on the mass of the leading pair (left) and of the sub-leading pair (right). The probability refers to a 125 GeV Higgs signal and to an integrated luminosity of 20 fb^{−1} at a cetre-of-mass energy of 7 TeV.

Figure 2: Dependency of the expected Z_{0} as described in the text on the mass of the leading pair. (right) dependendy of expected Z_{0} as described in the text on the transverse momenta of the third and fourth muons in the quadruplet. The probability refers to a 125 GeV Higgs signal and to an integrated luminosity of 20 fb^{−1} at a cetre-of-mass energy of 7 TeV.

128 In addition the impact of extending the muon acceptance by adding the SA and Calo-Tag muons as 129 defined in Section 3 has been investigated. The SA muons largely improves the efficiency in the region 130 at large pseudo-rapidities (2.5 < |η| < 2.7), where the Inner Detector tracking efficiency drops rapidely, 131 and the Calo-Tag muons in the pseudo-rapidity region |η| < 0.1 where there are not instrumented regions 132 of the Muon Spectrometer due to the passage of ATLAS services. By adding the SA and Calo-Tag muon 133 the signal efficiency, for a 125 GeV Higgs mass, is improved by about 7% in the otpimized kinematic 134 region. The details on the change of the selection, with respect to the published one, and the expected 135 signal significance are reported in Table 3 for an integrated luminosity of 20 fb^{−1} at a centre-of-mass 136 energy of 7 TeV for the four-muons channel alone and a Higgs signal with a mass of 125 GeV. 137 Usign the same statistical method the impact of other muon selection variables on signal sensitivity 138 was investigated. In particular the use of the calorimetric isolation with pileup corrections was showing 139 a decrease of sensitivity due to the increase of the reducible background. In this case it has been decided 140 to use the uncorrected calorimetric isolation that gives not only better sensitivity but also a sizable reduc-141 tion of the reducible background (conservative approach). Similar study confirmed the improvement in

Variable Optimal Cut  Published cut
M_{12}  50 GeV  15 GeV around Z mass
M_{34}  20 GeV  17.5 GeV
P_{t1}  20 GeV  20 GeV
P_{t2}  15 GeV  20 GeV
P_{t3}  10 GeV  7 GeV
P_{t4}  6 GeV  7 GeV

Table 2: Value of optimal cut position for a 125 GeV Higgs mass signal obtained form the procedure described in the text for the four muon final state. The M_{34} cut values refers to a M_{4μ} mass of 125 GeV.

Selection  Signal Irreduc. bkg. Reduc. bkg. Z_{0} Gauss. σ
Published without SA and calo-tag μ 2.25  1.26  0.08  1.64
Published with SA and calo-tag μ  2.40  1.37  0.10  1.69
Optimized with SA and calo-tag μ  3.66  1.92  0.20  2.14

Table 3: Number of signal and background events for the different selections from simulation assuming an integrated luminosity of 20 fb^{−1} at a centre-of-mass energy of 7 TeV for the four-muons channel alone and a Higgs signal with a mass of 125 GeV. No errors on the estimated background is considered here. In the last column the expected local signal significnace, P_{0}, expressed in terms of equivalent one-sided Gaussian sigma is reported.

142 sensitivity if the impact parameter significance cut is applied to all four muons of the quadruplet.

143 4.2 Optimization for channels with Electrons

144 Due to the different background composition a separate optimization study was performed for channels 145 with sub-leading electrons final states, i.e. the 4e and 2μ2e channels. The kinematic cuts in the quadru-146 plet selection were studied in details after fixing the minimum transverse momentum P_{te}> 7 GeV. As in 147 the previous subsection, the main variables studied were the closest mass to the Z one, M_{12}, the lowest 148 momentum of the electrons in the first pair, P_{t(1,2)}, and the transverse momentum of the lower energy 149 electrons P_{t3}, while keeping the electron with the lowest momentum P_{t4}> 7 GeV. The optimization strat-150 egy was to focus on the low-mass Higgs region where potential changes in the kinematic cuts can impact 151 both the signal and reducible backgrounds. 152 The observables used to quantify the acceptance to MC signal and background events of a given set 153 of cuts are the expected number of events for the integrated luminosity of 2011 (4.8 fb^{−1} ) and the global 154 and local significances estimated with Eq. (1). While the global significance looks at a wide range of 155 Higgs mass values (100-180 GeV), the local significance is computed in a limited range of ±6 GeV 156 (corresponding roughly to 2×FWHM of the experimental resolution) centered on a given Higgs mass 157 hypothesis. The optimization procedure, in parallel with the muon-based procedure, was to re-run the 158 standard analysis with fixed electron and muon identification and isolation event selections on signal and 159 background MC, while varying the targeted variables. Note that the reduction in Z+jets backgrounds is 160 accounted for with the implementation of the new Eratio cut for low energy electrons with P_{te}< 10 GeV. 161 See Section Section 3 for more details. 162 Figure 3 shows the expected number of events (left), the global (center) and Local (right) signifi-163 cances as a function of the different kinematic cuts for an integrated luminosity of 4.8 fb^{−1} . The top 164 row of Figure 3 refers to the M_{12}, the middle to the P_{t1} abd the bottom to the P_{t3} kinematic variables 165 respectively. 166 Figure 3 top clearly show that lowering the value of M_{12} with respect to the published analysis 167 increases the acceptance of the analysis to both signal and background events. Passed the lower bound 168 of M_{12} > 50 GeV, the local and global significances reach a stable value where no more significant 169 improvement is possible. For this cut value, the increase in the expected number of events is nearly 170 ∼70% for a Higgs signal at 120 GeV. 171 Figure 3 middle show that lowering the sub-leading P_{t2} gives a small improvement on the signifi-172 cances. If it is lowered to 15 GeV, with respect to the previous value of 20 GeV, the gain in the global 173 significance is at the level of ∼ 4% although the number of Zbb+jets events have increased for a 120-GeV 174 Higgs. 175 Figure 3 bottom show that further improvement can be reached by increasing the minimum P_{t} of 176 the third lepton to 10-12 GeV, with respect to the default lowest P_{te} > 7 GeV due to a better rejection 177 of Z+jets and tt̄backgrounds characterized by the negative differences in the number of events in Fig-178 ure Figure 4.2. 179 In light of the previous one-dimensional kinematic cut parameter scans, it appears reasonable given 180 the increases in the number of events and both local and global significances to improve the analysis 181 sensitivity at low Higgs mass with a different set of kinematic cuts compared to the published analysis. 182 The suggested new values for three dominant cuts were found to be M_{12} > 50 GeV, P_{t(2)} > 15 GeV, 183 and P_{t3}> 10 GeV. The latter exact values can be fine-adjusted to make the analysis of all sub-channels 184 consistent without great impact on the figures stated above.

Figure 3: Expected number of events (left) for an integrated luminosity of 4.8fb^{−1} , global (middle) and local (right) significances as a function of the lower bound of M_{12} (top row), the sub-leading P_{t2}found in the first lepton pair (middle row), and the P_{t3} of the third lepton (bottom row). The bottom part of each plot gives the relative difference, in percent, with respect to thepublished analysis.

185 5 Event Selection

May 9, 2012 – 09 : 24  DRAFT  12

186 6 Trigger Considerations

187 Contact : S. Rosati

188 7 Data based study of the lepton selection criteria

189 Contact : A. Schaffer

190 7.1 Efficiency for signal-like leptons

191 7.1.1 Z → ee 192 7.1.2 Z → μμ

193 7.2 Efficiency for background-like leptons

194 7.2.1 Z + e 195 7.2.2 Z + μ

196 8 Background Estimation

197 Contact: C. Anastopoulos

198 8.1 Data driven cross-checks of the tt̄background

199 8.2 Estimate of the Z + μμ background

200 8.2.1 Efficiency estimation for muons in Jets

201 The estimation of the reducible background Z(→ μμ) + μμ is important to determine it’s contribution 202 to the H → 4μ channel. An important step in this estimation process is to determine the efficiency of 203 the isolation cuts on the muons. In order to obtain a sample with a greater level of statistics we selected 204 events with a Z that decays into two muons plus an additional muon. The selection of the Z is performed 205 by requiring two opposite sign combined or tagged muons with p_{T 1}> 20 GeV and p_{T 2}> 10 GeV and 206 an invariant mass M_{Z} > 50 GeV. The isolation and impact parameter significance cuts are the same as 207 those of the baseline selection. The additional muon is allowed to be standard (combined or tagged), 208 standalone or calo. Each of the three categories has its own p_{T} and η cuts:

209 - additional combined or tagged muon: p_{T} > 5 GeV and |η| < 2.7

210 - additional standalone muon: p_{T} > 5 GeV and |η| > 2.5

211 - additional calo muon: p_{T} > 15 GeV and |η| < 0.1.

212 The observed number of events and the escpected number of background events for the different MC samples assuming the same integrated luminosity of the data are reported in Table 4. In Figure 4 the

Sample  Z + 1  Z + 1  Z + 1
combined or tagged μ standalone μ  calo μ
Events  Events  Events
ZZ  33 ± 0.16  4.2 ± 0.055 2.9 ± 0.046
Zbb  2539 ± 6.9  43 ± 0.89  5.4 ± 0.31
Zjets  7348 ± 66  165 ± 9.9  209 ± 11
tt̄  966 ± 5.3  8.4 ± 0.50  3.4 ± 0.32
WZ  195 ± 4.2  9.00 ± 0.89 5.3 ± 0.68
DY  4.3 ± 2.1  0  0
Total BKG  11085 ± 67  230 ± 10  226 ± 11
Data  12888 ± 114  230 ± 15  172 ± 13

Table 4: Observed number of events and expected number of background events for the different MC samples assuming the same integrated luminosity of the data.

213

214 reconstructed invariant mass of the Z is shown for each MC background and the data, in the case of 215 events with a Z and an additional combined or tagged muon. In the left plot the MC distributions are 216 weighted by the cross sections and normalized to the data luminosity. In the right plot the distributions 217 of the reconstructed invariant mass of the Z for the total MC background and the data are normalized to 218 1. The shape of the distributions is well described by the MC and the normalization agreement in terms 219 of number of events is better then 15% when the third muon is combined or tagged and even better in the 220 other two cases. 221 To estimate the rejection power of the calorimetric isolation a cut on the transverse energy inside a cone 222 with an opening angle of 0.20 around the additional muon divided by its transverse momentum has been Figure 4: Reconstructed invariant mass of the Z for both MC and data in the case of events with a Z plus an additional combined or tagged muon. In the left plot the MC distributions are weighted by the cross sections and normalized to the data luminosity. In the right plot the same distributions are normalized to 1.

223 applied. The distributions of the calorimetric isolation variable E_{T cone20}/p_{T} for the additional muon are 224 shown in Figure 5 for the case of a combined or tagged muon and in Figure 6 for a standalone (left plot) 225 and calo muon (right plot). The number of events with E_{T cone20}/p_{T}< 0.3 and the efficiency of this cut are 226 listed in Table 5 for all three cases. The normalization agreement is always better than the 17% level while 227 the efficiencies of this cut for MC and data are compatible. In Figure 5 the E_{T cone20}/p_{T} distributions for 228 the additional combined or tagged muon are weighted by the cross sections and normalized to the data 229 luminosity (left plot) and normalized to 1 (right plot). The normalized plot shows that the shape of the 230 data and MC distribution are in a quite good agreement. When the third muon is a standalone muon an 231 additional tighter cut has also been applied and the efficiency for the selection E_{T cone20}/p_{T} < 0.15 has

been estimated (see Table 6). Data and MC for both normalization and efficiency are in agreement.

Figure 5: Distributions of the calorimetric isolation variable E_{T cone20}/p_{T} for the additional muon in the case of a combined or tagged muon when MC backgrounds are normalized to the integrated luminosity of the data (left plot) and when MC and data are normalized to 1 (right plot).

232

233 A selection on the track isolation has been applied with a cut on the transverse momentum inside 234 a cone with an opening angle of 0.20 around the additional muon divided by its transverse momentum. 235 The distributions of the p_{T cone20}for the additional muon are shown in Figure 7 for a combined or tagged Figure 6: Distributions of the calorimetric isolation variable E_{T cone20}/p_{T} for the additional muon in the case of a standalone (left plot) and a calo muon (right plot).

Sample  Z + 1  Z + 1  Z + 1
combined or tagged μ  standalone μ  calo μ
Events Efficiency (%)  Events  Efficiency (%)  Events  Efficiency (%)
ZZ  32 ± 0.15  97.9 ± 2.5  4.11 ± 0.05  99 ± 5  2.87 ± 0.05  99 ± 4
Zbb  1331 ± 5  52.4 ± 0.7  26.7 ± 0.7  62 ± 6  4.0 ± 0.3  74 ± 16
Zjets  3232 ± 44  44.0 ± 0.4  123.4 ± 9  75 ± 3  25 ± 2  122 ± 8
tt̄  208 ± 3  21.5 ± 0.6  3.4 ± 0.3  41 ± 11  2.1 ± 0.2  61 ± 21
WZ  173 ± 4  88.3 ± 2.2  8.4 ± 0.9  93 ± 8  7 ± 2  93 ± 10
DY  4.3 ± 2.1  0  0
Total BKG 4980 ± 44  44.9 ± 0.3  166 ± 9  72 ± 3  136 ± 9  60 ± 3
Data  6025 ± 78  46.7 ± 0.3  167 ± 13  73 ± 3  90 ± 9  52 ± 3

Table 5: Observed number of events and expected number of background events for the different MC samples assuming the same integrated luminosity of the data and requiring E_{T cone20}/p_{T} < 0.3.

236 muon (left plot) and for a calo muon (right plot). The number of events with p_{T cone20}/p_{T} < 0.15 and the 237 efficiency of this cut are listed in Table 7 for the two cases of an additional combined or tagged muon 238 and an additional calo muon. The normalization agreement is always better than the 15% level while the 239 efficiencies for MC and data are compatible. 240 To estimate the rejection power of the impact parameter significance selection a cut on the impact 241 parameter significance has been applied. The distributions of the impact parameter significance d_{0}/σ(d_{0}) 242 for the additional muon are shown in Figure 8 for a combined or tagged muon (left plot) and for a calo 243 muon (right plot). The number of events with d_{0}/σ(d_{0}) < 3.5 and the efficiency of this cut are listed in 244 Table 8 for the case of a combined or tagged muon and for the case of a calo muon. The normalization 245 agreement is always better than the 15% level, while the efficiencies for MC and data are compatible. 246 The observed number of events, the expected number of background events, and the efficiencies 247 considering the combined effect of the two isolation cuts and the impact parameter significance cut for 248 the different MC backgrounds and the data in the case of a Z plus an additional muon are listed in Table 249 9 for a combined or tagged additional muon, Table 10 for a standalone muon and Table 11 for a calo 250 muon.

Sample  Z + 1

standalone μ

Events  Efficiency (%)
ZZ  4.04 ± 0.05  97 ± 8
Zbb  20.2 ± 0.6  47 ± 5
Zjets  104 ± 8  63 ± 3
tt̄  2.2 ± 0.2  26 ± 8
WZ  8.2 ± 0.8  91 ± 9
DY  0
Total BKG 139 ± 8  60 ± 2
Data  139 ± 12  60 ± 2

Table 6: Observed number of events and expected number of background events for the different MC samples assuming the same integrated luminosity of the data and requiring E_{T cone20}/p_{T} < 0.15.

Figure 7: Distribution of the transverse momentum inside a cone with an opening angle of 0.20 around the additional muon, p_{T cone20}, in the case of a combined or tagged additional muon (left plot) and a calo additional muon (right plot).

251 8.3 Estimate of the Z + ee background

252 8.4 Data driven cross-checks of the QCD background

253 8.5 Background shapes

Sample  Z + 1  Z + 1
combined or tagged μ  calo μ
Events Efficiency (%)  Events  Efficiency (%)
ZZ  32.1 ± 0.2  98 ± 3  2.86 ± 0.05  99 ± 6
Zbb  962 ± 4  37.9 ± 0.6  2.6 ± 0.2  48 ± 15
Zjets  2587 ± 39  35.2 ± 0.3  78 ± 7  37 ± 2
tt̄  137 ± 2.0  14.1 ± 0.4  0.8 ± 0.2  24 ± 11
WZ  4.8 ± 0.6  90 ± 12
DY  4 ± 2  86 ± 15  0
Total BKG 3893 ± 40  35.1 ± 0.3  89 ± 7  39 ± 2
Data  4593 ± 68  35.6 ± 0.3  65 ± 8  38 ± 2

Table 7: Observed number of events and expected number of background events for the different MC samples assuming the same integrated luminosity of the data and requiring p_{T cone20}/p_{T} < 0.15.

Figure 8: Distributions of the impact parameter significance d_{0}/σ(d_{0}) for an additional combined or tagged muon (left plot) and for a calo muon (right plot).

Sample  Z + 1  Z + 1
combined or tagged μ  calo μ
Events Efficiency (%)  Events  Efficiency (%)
ZZ  29.3 ± 0.2  89 ± 5  2.86 ± 0.05  99 ± 6
Zbb  1232 ± 5  48.5 ± 0.7  3.6 ± 0.2  66 ± 17
Zjets  6051 ± 60  82.4 ± 0.4  199 ± 11  95 ± 1
tt̄  410 ± 3  42 ± 1  2.2 ± 0.2  63 ± 21
WZ  186 ± 4  95 ± 2  5.2 ± 0.7  98 ± 6
DY  4 ± 2  100.0 ± 0.5  0
Total BKG 7913 ± 60  71.4 ± 0.4  213 ± 11  94 ± 2
Data  9282 ± 96  72.0 ± 0.3  163 ± 13  95 ± 2

Table 8: Observed number of events and expected number of signal and background events for the different MC samples assuming the same integrated luminosity of the data and requiring d_{0}/σ(d_{0}) < 3.5.

Sample  Z + 1

combined or tagged μ

Events Efficiency (%) Rejection (%)
ZZ  28.4 ± 0.1  86 ± 6  14 ± 2
Zbb  461 ± 3  18.2 ± 0.3  81.8 ± 0.7
Zjets  1654 ± 31  22.5 ± 0.2  77.5 ± 0.4
tt̄  59 ± 1  6.1 ± 0.2  93.9 ± 0.7
WZ  163 ± 4  84 ± 2  16 ± 1
DY  4 ± 2  86 ± 15  14 ± 6
Total BKG 2369 ± 32  21.4 ± 0.2  78.6 ± 0.3
Data  2935 ± 54  22.8 ± 0.2  77.2 ± 0.3

Table 9: Observed number of events and expected number of background events for the different MC samples assuming the same integrated luminosity of the data and requiring E_{T cone20}/p_{T} < 0.3, p_{T cone20}/p_{T} < 0.15, d_{0}/σ(d_{0}) < 3.5 in the case of a combined or tagged muon.

Sample  Z + 1

standalone μ

Events  Efficiency (%) Rejection (%)
ZZ  4.10 ± 0.05  98 ± 6  1.5 ± 0.7
Zbb  26.1 ± 0.7  60 ± 6  40 ± 5
Zjets  114 ± 8  69 ± 3  31 ± 2
tt̄  3.3 ± 0.3  40 ± 11  60 ± 13
WZ  8.4 ± 0.9  93 ± 8  7 ± 2
DY  0
Total BKG 156 ± 8  68 ± 3  32 ± 2
Data  159 ± 13  69 ± 3  31 ± 2

Table 10: Observed number of events and expected number of background events for the different MC samples assuming the same integrated luminosity of the data and requiring E_{T cone20}/p_{T} < 0.15, p_{T cone20}/p_{T} < 0.15, d_{0}/σ(d_{0}) < 3.5 in the case of a standalone muon.

Sample  Z + 1
calo μ
Events  Efficiency (%) Rejection (%)
ZZ  2.80 ± 0.05  97 ± 10  3 ± 2
Zbb  2.0 ± 0.2  38 ± 13  62 ± 17
Zjets  57 ± 6  27 ± 2  73 ± 3
tt̄  0.6 ± 0.1  19 ± 9  81 ± 19
WZ  4.6 ± 0.6  87 ± 14  13 ± 5
DY  0
Total BKG  67 ± 6  29 ± 2  71 ± 3
Data  48 ± 7  28 ± 2  72 ± 3

Table 11: Observed number of events and expected number of background events for the different MC samples assuming the same integrated luminosity of the data and requiring E_{T cone20}/p_{T} < 0.3, p_{T cone20}/p_{T} < 0.15, d_{0}/σ(d_{0}) < 3.5 in the case of a standalon muon.

254 9 Results of Event Selection

May 9, 2012 – 09 : 24  DRAFT  21

255 10 Systematic Uncertainties

256 Contact : J. Hoffman

May 9, 2012 – 09 : 24  DRAFT  22

257 11 Exlusion limits and p-values

258 Contact : L. Flores

259 12 Conclusions

24

Appendices^{260}